# DRWKV: Focusing on Object Edges for Low-Light Image Enhancement

### v2

@@ VII-B1 Evolving Spatial Mix
$W_{(*)}$ denotes the weight. To ensure numerical stability, prevent training divergence, and accelerate training convergence, we introduce the LN layer. The total output of the Evolving Spatial Mix block is as follows:
@@ VII-B2 Channel Mix
$\textit{where }V_{c}=SR((O_{s}+(1-\mu_{(k)})O_{s}^{\dagger})W_{K})W_{V}$ (28)
VII-C Bilateral Spectrum Aligner
@@
VII-C1 Spectral Alignment Enhancer (SAE)
@@ VII-C2 Cross Attention (CA)
VII-C3 Scharr Edge Enhancement (SEE)
@@
TABLE V: User study on challenging low-light image enhancement (mean opinion score).
Metric
RetinexNet Zero-DCE RRDNet RetinexFormer RetinexMamba MambaIR DRWKV
Overall Rating 4.08 3.94 3.72 3.97 3.92 4.67 4.87
Local Details 3.95 4.04 3.72 4.03 3.92 3.96 4.85
Color Restoration 3.89 3.76 3.67 4.05 4.03 4.64 4.90
Total 3.97 3.91 3.70 4.02 3.96 4.42 4.87
@@
Finally, to achieve adaptive adjustment of feature dimensions and deep fusion of semantic information, $1\times 1$ convolution and $3\times 3$ convolution are used for dimension reduction and recovery, resulting in the final output $E\in\mathbb{R}^{C\times H\times W}$.
@@
$E=Conv_{3\times 3}(SiLU(Conv_{1\times 1}(E_{Fusion})))$ (47)
@@
VIII Extended experiment
VIII-A User Subjective Evaluation Experiment
To rigorously assess the perceptual quality of enhancement algorithms under complex real world conditions we designed and conducted a user subjective evaluation experiment. Ten low light images were randomly selected from the publicly available LSRW Huawei benchmark and processed by seven enhancement models: RetinexNet, Zero-DCE, RRDNet, RetinexFormer, RetinexMamba, MambaIR and the proposed DRWKV to generate corresponding enhanced results. One hundred participants spanning diverse ages and professional backgrounds were then recruited. In an independent double blind protocol participants assigned scores on a five point Likert scale from one worst to five best along three key perceptual dimensions overall visual quality local detail restoration and global color fidelity. The mean opinion scores are reported in Table V. Across all three dimensions the proposed DRWKV model achieved the highest ratings indicating superior perceptual enhancement performance from the human visual perspective.
VIII-B Low-light Object Tracking Experiment

### v3

@@ 7.2.1 Evolving Scanning Spatial Mix
$W_{(*)}$ denotes the weight. To ensure numerical stability, prevent training divergence, and accelerate training convergence, we introduce the LN layer. The total output of the Evolving Scanning Spatial Mix block is as follows:
@@ 7.2.2 Channel Mix
$\textit{where }V_{c}=SquaredReLU(K_{c})$ (30)
7.3 Bilateral Spectrum Aligner Block
@@
7.3.1 Spectral Alignment Enhancer (SAE)
@@ 7.3.2 Cross Attention (CA)
7.3.3 Scharr Edge Enhancement (SEE)
@@
7.4 User Subjective Evaluation Experiment
Figure 9: User subjective evaluation experiment.
@@ 7.4 User Subjective Evaluation Experiment
Finally, to achieve adaptive adjustment of feature dimensions and deep fusion of semantic information, $1\times 1$ convolution and $3\times 3$ convolution are used for dimension reduction and recovery, resulting in the final output $SEE_{out}\in\mathbb{R}^{C\times H\times W}$.
@@
$W_{x}=\begin{bmatrix}-3&0&3\\ -10&0&10\\ -3&0&3\end{bmatrix},W_{y}=\begin{bmatrix}-3&-10&-3\\ 0&0&0\\ 3&10&3\end{bmatrix}$ (47)
@@
$SEE_{out}=Conv_{3\times 3}(SiLU(Conv_{1\times 1}(E_{Fusion})))$ (50)
@@
8 The input of the Global Edge Retinex theory
In the main text, we elaborate and derive the proposed GER, and present the input structure in the main text figures: edge features $E$, artifacts $S$, and processed global illumination $L_{ill}$. As the artifacts $S$ and processed global illumination $L_{ill}$ have been thoroughly explained, in the appendix, we discuss the extraction and integration of edge features $E$.
To reduce the computational load and complexity of the model itself, we further utilize the SEE based on the output $I_{DRWKV}\in\mathbb{R}^{3\times H\times W}$ of the DRWKV model, thereby obtaining the integration of input edge features $E$ suitable for Global Edge Retinex theory.
The specific calculation process can be expressed as follows:
$E=SEE(I_{DRWKV})$ (51)
9 Extended experiment
9.1 User Subjective Evaluation Experiment
To rigorously assess the perceptual quality of enhancement algorithms under complex real world conditions we designed and conducted a user subjective evaluation experiment. Ten low light images were randomly selected from the publicly available LSRW Huawei benchmark and processed by seven enhancement models: RetinexNet, Zero-DCE, RRDNet, RetinexFormer, RetinexMamba, MambaIR and the proposed DRWKV to generate corresponding enhanced results. One hundred participants spanning diverse ages and professional backgrounds were then recruited. In an independent double blind protocol participants assigned scores on a five point Likert scale from one worst to five best along three key perceptual dimensions overall visual quality local detail restoration and global color fidelity. The mean opinion scores are reported in Figure 9. Across all three dimensions the proposed DRWKV model achieved the highest ratings indicating superior perceptual enhancement performance from the human visual perspective.
9.2 Low-light Object Tracking Experiment
@@
9.3 $\alpha$, $\beta$ and $\gamma$ Parameter Detail Experiment
To obtain the optimal configuration of the three parameters ($\alpha$, $\beta$ and $\gamma$), we designed a series of parameter detail experiments. For the dataset, we adopted the LOLv2-Real dataset. During the implementation of the parameter detail experiments, we only adjusted one of the three parameters ($\alpha$, $\beta$ and $\gamma$) while keeping the other two unchanged. The experimental results are presented in Table 5, Table 6, and Table 7. These results indicate that the optimal parameter configuration for DRWKV is $\alpha=0.2$, $\beta=-0.05$, and $\gamma=0.2$.
Table 5: The impact of different $\alpha$ values on the performance of the LOLv2-Real dataset.
  $\alpha$ 0.05 0.10 0.15 0.20 0.25 0.30 0.35
PSNR↑ 23.56 23.89 24.02 24.12 23.84 23.97 23.94
SSIM↑ 0.798 0.815 0.814 0.832 0.792 0.821 0.814
NIQE↓ 4.012 3.967 3.941 3.926 3.951 3.989 4.173
Table 6: The impact of different $\beta$ values on the performance of the LOLv2-Real dataset.
   $\beta$ -0.100 -0.075 -0.050 -0.025 0
PSNR↑ 23.72 23.98 24.12 23.86 23.61
SSIM↑ 0.803 0.821 0.832 0.817 0.804
NIQE↓ 3.995 3.948 3.926 3.973 4.003
Table 7: The impact of different $\gamma$ values on the performance of the LOLv2-Real dataset.
  $\gamma$ 0.05 0.10 0.15 0.20 0.25 0.30 0.35
PSNR↑ 22.74 23.65 23.91 24.12 23.99 23.78 23.84
SSIM↑ 0.804 0.801 0.823 0.832 0.826 0.812 0.814
NIQE↓ 4.112 4.023 3.957 3.926 3.962 3.998 4.472